OpenStudy (anonymous):

a straight road goes nearby a residential building, 50 m from the road . How much farther away would the house be placed so that the sound intensity level would have been 5 dB less. The traffic is unchanged

OpenStudy (anonymous):

I have to use these equations \[I=\frac{ P }{ 4\pi r }\] \[\beta=10*\log(\frac{ I }{ I_0 })\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

I did get a tip from my professor that I should count as the intensity decreases with 1/r, where r is the distance from the road...

OpenStudy (michele_laino):

we have to find the value I_1 such that: \[{\beta _1} = 10 \times \log \left( {\frac{{{I_1}}}{{{I_0}}}} \right) = \beta - 5 = 10 \times \log \left( {\frac{I}{{{I_0}}}} \right) - 5\]

OpenStudy (michele_laino):

or: \[\Large 10 \times \log \left( {\frac{I}{{{I_1}}}} \right) = 5\] where: \[\Large I\] is the initial intensity. So we have to know what is P, in your previous formula

OpenStudy (anonymous):

yea.. but how?

OpenStudy (michele_laino):

do you know the power P?

OpenStudy (anonymous):

No, I no nothing else than what Ive written in the question:S

OpenStudy (michele_laino):

please wait, I try to solve your question

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

I think that we can use this formula: \[\Large \frac{I}{{{I_1}}} = {\left( {\frac{r}{{{r_1}}}} \right)^2}\] where: \[r = 50\;meters\] and \[{{r_1}}\] is the requested distance I think that the right formula is: \[\Large I = \frac{P}{{4\pi {r^2}}}\]

OpenStudy (michele_laino):

so, what is : \[\large \frac{I}{{{I_1}}} = ...?\]

OpenStudy (anonymous):

what does \[\frac{ I }{ I_1 } =(\frac{ r }{ r_1 })^2\] comes from?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

since the power is the same, so we can write: \[\Large 4\pi {r^2}I = 4\pi r_1^2{I_1}\]

OpenStudy (michele_laino):

namely the power of the emitted sound doesn't change, since the traffic is unchanged

OpenStudy (anonymous):

I dont understand the algebra how you get it to \[\frac{ I }{ I_1} =(\frac{ r }{ r_1})^2\]

OpenStudy (anonymous):

\[\frac{ I }{ I_1 }=(\frac{ r_1 }{ r })^2\]

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

I get it to the last I wrote

OpenStudy (michele_laino):

yes! Sorry I have made an error, the right formula is: \[\Large \frac{{{I_1}}}{I} = {\left( {\frac{r}{{{r_1}}}} \right)^2}\]

OpenStudy (anonymous):

So \[\frac{ I_1 }{ (-5) }=(\frac{ 50 }{ r })^2\] @Michele_Laino

OpenStudy (michele_laino):

no, we have to use the condition above, so we get: \[\Large \frac{I}{{{I_1}}} = \sqrt {10} = {\left( {\frac{{{r_1}}}{{50}}} \right)^2}\] @pate16

OpenStudy (anonymous):

what, where does\[\sqrt{10}\] comes from?

OpenStudy (anonymous):

im sorry for the questions.. @Michele_Laino

OpenStudy (michele_laino):

no worries :) it comes from the condition that the new intensity has to be 5 dB less than the initial intensity

OpenStudy (anonymous):

and now you did write \[\frac{ r_1 }{ r }\] again?

OpenStudy (michele_laino):

yes! and r=50 meters

OpenStudy (anonymous):

yah but you corrected it and said \[\frac{ r }{ r_1 }\] ?

OpenStudy (michele_laino):

as you well commented, we have: \[\frac{I}{{{I_1}}} = {\left( {\frac{{{r_1}}}{r}} \right)^2}\]

OpenStudy (anonymous):

sorry sorry!!!! but can you explain the squareroot 10 again?

OpenStudy (michele_laino):

so, substituting r = 50 meters, we have: \[\frac{I}{{{I_1}}} = {\left( {\frac{{{r_1}}}{{50}}} \right)^2}\]

OpenStudy (michele_laino):

your problem requests, that the new intensity I, has to be such that the subsequent quantity: \[{\beta _1} = 10 \times \log \left( {\frac{{{I_1}}}{{{I_0}}}} \right)\] is 5 less than this quantity: \[\beta = 10 \times \log \left( {\frac{I}{{{I_0}}}} \right)\] where: \[I\] is the initial intensity, whereas \[{{I_1}}\] is the new intensity, and \[{{I_0}}\] is a refernce intensity, whose value is a standard value

OpenStudy (michele_laino):

so we can write: \[\large 10 \times \log \left( {\frac{{{I_1}}}{{{I_0}}}} \right) = 10 \times \log \left( {\frac{I}{{{I_0}}}} \right) - 5\]

OpenStudy (michele_laino):

or, using the properties of logarithm: \[\large \begin{gathered} 10 \times \log \left( {\frac{I}{{{I_0}}}} \right) - 10 \times \log \left( {\frac{{{I_1}}}{{{I_0}}}} \right) = 5 \hfill \\ \hfill \\ \log \left( {\frac{I}{{{I_1}}}} \right) = \frac{1}{2} \hfill \\ \hfill \\ \frac{I}{{{I_1}}} = {10^{1/2}} = \sqrt {10} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

yeah im with you now

OpenStudy (anonymous):

\[\sqrt{10}=(\frac{ r1 }{ 50 })^2\]

OpenStudy (anonymous):

solve for r?

OpenStudy (anonymous):

@Michele_Laino 88.91 m?

OpenStudy (michele_laino):

that's right! @pate16

OpenStudy (anonymous):

damn it your good michele!!

OpenStudy (michele_laino):

thanks! :)

OpenStudy (anonymous):

Thanks alot!!!!